OpenStudy (anonymous):

PLEASE HELP ME FAST At the school cafeteria, the cost of one milk is $0.50. The total cost of 5 turkey sandwiches and 4 milks is the same as the total cost of 4 turkey sandwiches and 8 milks. a. Write an equation that can be used to find the cost of a turkey sandwich. Explain what each expression or value represents in the equation. b. Use the equation you wrote in part a. to find the cost of a turkey sandwich. Explain how you solved the equation.

OpenStudy (texaschic101):

t = turkey sandwiches m = milk 5t + 4m = 4t + 8m 5t - 4t = 8m - 4m t = 4m m = .50 sub .50 in for m in your answer above t = 4m t = 4(.50) t = 2.00 lets check.... 5t + 4m = 4t + 8m 5(2) + 4(.5) = 4(2) + 8(.5) 10 + 2 = 8 + 4 12 = 12 (correct) ANSWER : turkey sandwiches are $2.00